Shutdown nanospec2 SNF 2011-06-24 16:33:45: wavelength counter is stuck

vilanova at snf.stanford.edu vilanova at snf.stanford.edu
Thu Jun 30 15:00:59 PDT 2011


Vendor Replaced Counter and Nylon Gear , HUb, Aligned Head In dark Room tested Counter Repeatable , Tested Ok. Good to use. 




More information about the nanospec2-pcs mailing list